OpenStudy (anonymous):

Solve the equation one-half times x plus three-fourths equals one-fourth for x. x equals one-fourth x equals negative one-fourth x equals 1 x equals negative 1

OpenStudy (anonymous):

okay so ....lets put those words in to numbers first :)

OpenStudy (anonymous):

kkk did u read my message

OpenStudy (anonymous):

solve: 1/2x+3/4=1/4x x=1/4 x=-1/4 x=1 x=-1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

okay well first lets get the x's to one side... 1/2x+3/4=1/4x 2/4x+3/4=1/4x 3/4=-1/4x

OpenStudy (anonymous):

did u subtract it and got 1/10

OpenStudy (anonymous):

umm...nope

OpenStudy (anonymous):

den wat u do

OpenStudy (anonymous):

wat was the first step that u didd

OpenStudy (anonymous):

message me

OpenStudy (anonymous):

you have to make all the denominators the same

OpenStudy (anonymous):

and then solve 1/2x + 3/4 = 1/4x 2/4x + 3/4 = 1/4x 3/4 = -1/4x

OpenStudy (anonymous):

kk thx so much would u help mee wit mor ques plz

OpenStudy (anonymous):

now what is x??

OpenStudy (anonymous):

i leav in 5 minutes so ill help you with what i can

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

quickly the question?? :)

OpenStudy (anonymous):

forget it idk knoe how to u cant copy and paste slopes

OpenStudy (anonymous):

but thx by the way

OpenStudy (anonymous):

biiii thhxxxx